# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

### v5

@@ Abstract
Despite major advances in machine learning, current artificial intelligence systems continue to fall short of human-like general intelligence. While large language and reasoning models can generate fluent and coherent outputs, they lack the deep understanding and adaptive reasoning that characterize truly general intelligence. Existing evaluation frameworks, which are centered on broad language or perception tasks, fail to capture generality at its core and offer no guidance. The artificial general intelligence testbed (AGITB) is a novel and freely available benchmarking suite comprising twelve fully automatable tests designed to evaluate low-level cognitive precursors through binary signal prediction. AGITB requires models to forecast temporal sequences without pretraining, symbolic manipulation, or semantic grounding. The framework isolates core computational invariants—such as determinism, sensitivity, and generalization—that align with principles of biological information processing. Engineered to resist brute-force and memorization-based approaches, AGITB presumes no prior knowledge and demands learning from first principles. While humans pass all tests, no current AI system has met the full AGITB criteria, underscoring its potential as a rigorous, interpretable, and actionable benchmark for guiding and evaluating progress toward artificial general intelligence. A reference implementation of AGITB is available on GitHub.
@@ 1 Introduction
The progress in AI has led to growing speculation that AGI is near. However, such claims remain difficult to substantiate without a rigorous and informative metric. Assessing progress toward AGI requires more than specialised benchmarks or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing essential cognitive capabilities.
@@
AGITB draws inspiration from the ladder to human-comparable intelligence introduced by Šprogar (2018), extending its conceptual foundations into a unified testbed comprising twelve automatable, domain-agnostic tests. While AGITB does not aim to assess consciousness or semantic comprehension, it provides a principled mechanism for distinguishing narrow AI systems from those exhibiting generalizable, adaptive behavior. A comparative analysis with the Abstraction and reasoning corpus (ARC) introduced by Chollet (2019) is included to contextualize AGITB within the broader landscape of AGI evaluation.
@@ 2 Background
2.1 A Test That Humans Pass but Machines Fail
@@
AGITB thus adopts a more elemental approach. Rather than judging intelligence by symbolic interpretation, it evaluates whether a system can detect, learn, and generalize patterns in raw binary signals independently of semantics or prior training. Rooted in the view—advanced by Hawkins and Blakeslee (2004)—that intelligence is fundamentally about extracting structure from data to make predictions, AGITB defines twelve core tests that conceptualize intelligence as a signal-based prediction. This approach aligns more closely with how biological intelligence appears to function at the cortical level—processing time-sensitive sensory data, not disembodied symbols.
@@ 3 Artificial General Intelligence Testbed
The primary objective of the artificial general intelligence testbed is to support the development and evaluation of AGI by defining a clear set of requirements that a model must meet to qualify as generally intelligent. A single model must fulfill the specific tasks defined in each test to validly claim success on the benchmark.
@@ 3.1 Components
Figure 1: (a) Example of a 2×32$\times$ 32 × 3-bit block of visual information combined with four bits of auditory data to form a single 10-bit input sample.
@@ 3.2 Operation
Figure 2: Adaptation to a sequence proceeds iteratively in discrete time steps. At the previous step ($t-1$), the model predicted that the current input would be $p^{*}_{t}$; this prediction was incorrect in the first bit of the auditory data. Based on the actual input $p_{t}$, the model then generates a new prediction $p^{*}_{t+1}$, which is incorrect in the fifth bit of visual data for the sequence shown in Figure 1b.
@@
To execute AGITB, users must specify one parameter: the pattern period. This value determines the length of the repeating temporal sequence in the input stream that the cortex is expected to recognize and adapt to. Longer pattern periods increase the temporal complexity of the task, requiring more sophisticated memory and generalization abilities. However, when combined with high-dimensional and wide inputs, excessively long periods may surpass the model’s learning capacity. As such, the pattern period should be selected in conjunction with input size to balance task difficulty with the model’s capabilities, ensuring that each test remains feasible and diagnostically informative.
3.3 The 12 Essential Tests
@@
Because the cortex updates its state based on cumulative history, the order of inputs critically affects learning and adaptation. Recognizing and exploiting temporal structure is thus a defining feature of intelligent systems.
Test 6 – Refractory Period
@@
Test 7 – Temporal flexibility
Assertion: The cortex must be able to adapt to temporal patterns with the user-specified period.
Assertion: The cortex must also be capable of adapting to temporal patterns with periods longer than the user-specified value.
This test evaluates a model’s capacity to generalize across temporal scales. Unlike rigid pattern-matching systems, intelligent models should flexibly detect and respond to recurring structures, regardless of their exact periodicity. The purpose of this test is to identify models that are overly specialized to a single temporal period rather than capable of robust generalization across multiple timescales.
Test 8 – Stagnation
@@
The structure of the input sequence affects how quickly a model can adapt. Simpler or more regular patterns typically lead to faster convergence, while complex or noisy inputs demand longer adjustment periods.
Test 10 – Salience
@@ 3.4 Memorization
[…]
@@
Consider an illustrative example of a task with 3-bit inputs and a pattern period of $n=5$. The total number of unconstrained sequences is $|S|=2^{15}=32.768$. Applying the refractory constraint reduces the valid set to $|S\textquoteright|=13^{3}=2.197$ sequences. When the cyclic condition is added, the space shrinks further to $|S^{\prime\prime}|=11^{3}=1331$.
Even with these biologically grounded constraints, the size of the valid sequence space grows exponentially concerning both sequence length and input dimensionality. Given that real-world sensory inputs may span millions of bits, strategies based on brute-force memorization are computationally intractable. The exponential complexity highlights the necessity for models that rely on generalizable, pattern-based learning mechanisms rather than rote memorization.
@@ 3.5.1 Human Performance
The inability to compare internal cortical states makes it impossible to directly verify AGITB tests in humans. However, because the cortical architecture naturally supports low-level binary signal processing and the tests align with cognitive expectations, we can assume that humans inherently satisfy most of the 12 tests. Tests 1 (Genesis) and 6 (Refractory Period), however, require additional argumentation.
Due to prior experience and cognitive bias, an adult human may appear to “fail” the first test (Genesis), since their cortex is no longer in a truly unconditioned state and may produce a non-empty prediction. In contrast, a fetal cortex—with no prior synaptic structure—receiving its first-ever input would satisfy the condition of true neutrality by predicting no spikes.
AGITB includes the refractory period requirement (Test 6) because it reflects biological reality; however, it does not eliminate the possibility of a deviating solution. Such a solution could feign compliance with the requirement without genuinely supporting it, whereas the converse is not possible.
Additionally, the more complex tests have counterparts at the cognitive level, where their effects can be observed through conscious reasoning and introspection. Temporal flexibility (Test 7), for example, is simple for humans as we have no problems recognizing temporal patterns of varying durations. The property of stagnation (Test 8) is a consequence of a (human) cortex being physically limited in its capacity to store knowledge. The test checks for the first sign of dementia—a moment when the cortex becomes saturated and forgets the recent event. Because humans learn different topics at varying speeds, the rationale for the entropy (Test 9) and the salience requirement (Test 10) is self-evident. The inability to directly inspect human mental states is captured in the unobservability requirement (Test 11). Finally, the generalization test (Test 12) requires what humans naturally excel at: generalizing when faced with new but distorted input.
@@ 3.5.4 Large Language Model Performance
Because LLM is essentially an ANN, it fails the genesis test. This failure immediately indicates that LLMs cannot learn entirely on their own and are inherently biased toward the language and data they were trained on. Although the internal state of an LLM is readily accessible and could, in principle, be used to perform other tests, such evaluations are meaningless until the first requirement is satisfied.
Consequently, LLMs cannot acquire new knowledge—let alone learn a new language—solely from prompts, nor can they develop the kind of grounded understanding of the world that humans possess. This implication also applies to large reasoning models, which are themselves built upon LLM architectures.
@@ 3.6 Remarks
While individual AGITB tests may seem trivial in isolation, the central challenge lies in designing a universal AGI architecture capable of satisfying all required tasks across the full suite within a cohesive and unified framework. As a recently proposed benchmark, AGITB has not yet been passed by any symbolic system, and it remains an open question whether such a solution is even theoretically possible. If a purely symbolic system were to succeed, it would suggest that AGITB fails to effectively differentiate narrow AI from genuinely general intelligence, given that symbolic systems—while proficient in syntactic manipulation—arguably lack true semantic understanding, as discussed by Searle (1980). Until such a system is demonstrated, AGITB’s assertions may be regarded as necessary (though not sufficient) conditions for identifying genuine, domain-independent intelligence. Although passing AGITB does not equate to achieving full AGI—since it does not, for instance, assess social reasoning or natural language fluency—failure to meet its conditions likely indicates a deficiency in foundational general cognitive capabilities.
@@ 4 AGITB and ARC: A Comparison
Task spatial dimensions 222 $\infty$

### v6

@@ Abstract
Despite major advances in machine learning, current artificial intelligence systems continue to fall short of human-like general intelligence. Existing evaluation frameworks, which are centered on language or perception tasks, fail to capture generality at its core and offer no guidance. The Artificial General Intelligence Testbed (AGITB) is a novel, freely available benchmarking suite consisting of thirteen core requirements, twelve of which are implemented as fully automatable tests designed to assess low-level cognitive precursors through binary signal prediction. AGITB requires models to forecast temporal sequences without pretraining, symbolic manipulation, or semantic grounding. The framework isolates core computational invariants—such as determinism, sensitivity, and generalization—that align with principles of biological information processing. Engineered to resist brute-force and memorization-based approaches, AGITB presumes no prior knowledge and demands learning from first principles. While humans pass all tests, no current AI system has met the full AGITB criteria, underscoring its potential as a rigorous, interpretable, and actionable benchmark for guiding and evaluating progress toward artificial general intelligence. A reference implementation of AGITB is available on GitHub.
@@ 1 Introduction
The progress in AI has led to growing speculation that AGI is near. However, such claims remain difficult to substantiate without a rigorous and informative metric. Assessing progress toward AGI requires more than specialized benchmarks or qualitative impressions; it necessitates principled, general-purpose metrics capable of systematically capturing and comparing essential cognitive capabilities.
@@
AGITB draws inspiration from the ladder to human-comparable intelligence introduced by Šprogar (2018), extending its conceptual foundations into a unified testbed comprising thirteen core requirements—twelve of which are implemented as fully automatable, domain-agnostic tests. While AGITB does not aim to assess consciousness or semantic comprehension, it provides a principled mechanism for distinguishing narrow AI systems from those exhibiting generalizable, adaptive behavior. A comparative analysis with the Abstraction and reasoning corpus (ARC) introduced by Chollet (2019) is included to contextualize AGITB within the broader landscape of AGI evaluation.
@@ 2 Background
2.1 A Benchmark That Only Humans and AGI Can Meet
@@
AGITB thus adopts a more elemental approach. Rather than judging intelligence by symbolic interpretation, it evaluates whether a system can detect, learn, and generalize patterns in raw binary signals independently of semantics or prior training. Rooted in the view—advanced by Hawkins and Blakeslee (2004)—that intelligence is fundamentally about extracting structure from data to enable predictions, AGITB operates at the level of signal-based forecasting. This approach aligns more closely with how biological intelligence appears to function at the cortical level—processing time-sensitive sensory data, not disembodied symbols.
@@ 3 Artificial General Intelligence Testbed
The primary objective of the artificial general intelligence testbed is to support the development and evaluation of AGI by defining a clear set of requirements that a model must meet to qualify as generally intelligent. A model must satisfy all requirements in order to validly claim success on the benchmark.
@@ 3.1 Components
Figure 1: (a) Example of a $2\times 3$-bit block of visual information combined with four bits of auditory data to form a single 10-bit input sample.
@@ 3.2 Operation
Figure 2: Adaptation to a sequence proceeds iteratively in discrete time steps. At the previous step (t1t-1), the model predicted that the current input would be $p^{*}_{t}$; this prediction was incorrect in the first bit of the auditory data. Based on the actual input $p_{t}$, the model then generates a new prediction $p^{*}_{t+1}$, which is incorrect in the fifth bit of visual data for the sequence shown in Figure 1b.
@@
To execute AGITB, users must specify one parameter: the temporal pattern length. This value determines the period of the repeating temporal sequence in the input stream that the cortex is expected to recognize and adapt to. Longer periods increase the temporal complexity of the task, requiring more sophisticated memory and generalization abilities. However, when combined with high-dimensional and wide inputs, excessively long periods may surpass the model’s learning capacity. As such, the temporal pattern length should be selected in conjunction with input size to balance task difficulty with the model’s capabilities, ensuring that each test remains feasible and diagnostically informative.
3.3 The 13 Essential Requirements
@@
Because the cortex updates its state based on cumulative history, the order of inputs critically shapes learning and adaptation. The ability to recognize and exploit temporal structure is therefore a defining property of intelligent systems.
However, sensitivity to input order alone is not sufficient. Since temporal coding makes time an intrinsic property of neural data, Requirement 6 is designed to rule out models that perform equally well on temporally invariant data, as such models fail to exploit temporal information.
Requirement 6 – Refractory Period
@@
Requirement 7 – Adaptability
Assertion: The cortex must be able to adapt to a temporal pattern with the user-specified period.
Assertion: The cortex must also be capable of adapting to a temporal pattern with a period longer than the user-specified value.
This test evaluates the model’s capacity to learn across temporal scales. Unlike rigid pattern-matching systems, intelligent models should flexibly detect and respond to recurring structures, regardless of their exact periodicity. The purpose of this test is to exclude models that cannot adapt to temporal patterns of different lengths.
Requirement 8 – Stagnation
@@
The structure and regularity of an input sequence affect the rate at which a model adapts, where adaptation time is defined as the number of iterations required to correctly predict the entire temporal pattern. Simpler or more predictable sequences typically result in faster convergence, while irregular or noisy inputs require greater exposure before the model can reliably learn and reproduce them.
Requirement 10 – Context sensitivity
@@
Requirement 13 – Bounded Prediction Latency
A cortex model must provide low-latency signal-level prediction, which is essential for functional viability. Biological brains achieve approximately constant reaction times through massive parallelism, such that per-input latency remains bounded and stable under expected workloads.
Unfortunately, no universal automatic test on serial hardware can verify this property across all possible cortex implementations within reasonable time. Instead, model authors must ensure that their architecture admits efficient parallel execution with bounded latency. In this respect, artificial neural networks—with their naturally parallelizable structure—offer a practical advantage.
@@ 3.4 Memorization
AGITB imposes a biologically inspired constraint: the refractory period, which prohibits any neuron (bit) from firing in consecutive time steps. This constraint significantly narrows the set of valid input sequences. The number of distinct length-$n$ temporal input sequences of $k$ bits, with the rule that a 1 never survives to the next time step, is $|S\textquoteright|=(F_{n+2})^{k}$, where $F_{i}$ denotes the $i$-th Fibonacci number, with $F_{0}=0$.
@@
Consider an illustrative example involving the learning of a temporal sequence consisting of $n=5$ input steps, each represented by a 10-bit binary vector. The total number of possible sequences is $|S|=2^{50}$. Introducing the refractory constraint significantly reduces this space to $|S\textquoteright|=(F_{7})^{10}=13^{10}$ valid sequences. Adding the cyclic condition further restricts the space to $|S^{\prime\prime}|=(L_{5})^{10}=11^{10}$ sequences.
Even with these biologically grounded constraints, the size of the valid sequence space grows exponentially concerning both sequence length and input dimensionality. Given that real-world sensory inputs can span tens of thousands of bits, brute-force memorization strategies become computationally intractable. This exponential complexity underscores the need for models that employ generalizable, pattern-based learning mechanisms rather than rote memorization.
@@ 3.5.1 Human Performance
The inability to compare internal cortical states makes it impossible to directly verify AGITB tests in humans. However, because the cortical architecture naturally supports low-level binary signal processing and the tests align with cognitive expectations, we can assume that humans inherently satisfy most requirements. Demand 1 (Genesis) and 6 (Refractory Period), however, require additional argumentation.
Due to prior experience and cognitive bias, an adult human may appear to “fail” the first prerequisite (Genesis), since their cortex is no longer in a truly unconditioned state and may produce a non-empty prediction. In contrast, a fetal cortex—with no prior synaptic structure—receiving its first-ever input would satisfy the condition of true neutrality by predicting no spikes.
AGITB includes the refractory period requirement (6) because it reflects biological reality; however, it does not eliminate the possibility of a deviating solution. Such a solution could feign compliance with the requirement without genuinely supporting it, whereas the converse is not possible.
Additionally, the more complex tests have counterparts at the cognitive level, where their effects can be observed through conscious reasoning and introspection. Temporal flexibility (Requirement 7), for example, is simple for humans as we have no problems recognizing temporal patterns of varying durations. The property of stagnation (Requirement 8) is a consequence of a (human) cortex being physically limited in its capacity to store knowledge. The test checks for the first sign of dementia—a moment when the cortex becomes saturated and forgets the recent event. Because humans learn different topics at varying speeds, the rationale for the content sensitivity (Requirement 9) and the context sensitivity (Requirement 10) is self-evident. The inability to directly inspect human mental states is captured in the unobservability requirement (Requirement 11). The generalization test (Requirement 12) requires what humans naturally excel at: generalizing when faced with new but distorted input. Finally, bounded prediction latency (Requirement 13) is a hallmark of the human brain.
@@ 3.5.4 Large Language Model Performance
Because an LLM is essentially an ANN, it fails the genesis test. This failure immediately indicates that LLMs cannot learn entirely from scratch and are inherently biased toward the language and data on which they were trained. Although an LLM’s internal state is readily accessible—and could, unlike that of a human, be inspected for further testing—any such evaluations are moot until the first requirement is satisfied.
If LLMs do not qualify as AGI, could they nevertheless simulate one? For example, what if an LLM were prompted to imitate a cortex that receives binary inputs, as specified by AGITB? A structured example of such a prompt is provided in the Appendix. However, before evaluating this type of simulation, we must revisit the sensitivity (Requirement 4), which stipulates that a model must remain sensitive to any single input, even after processing an arbitrarily long sequence of subsequent inputs.
This requirement is fundamentally incompatible with the architecture of LLMs, which operate within a fixed context window. Once this window is exceeded, earlier tokens are discarded or compressed, causing the system to forget or attenuate the long-term dependencies essential to satisfying the test.
More broadly, LLMs cannot acquire new knowledge—or learn a new language—purely from prompts. Nor can they develop the kind of grounded, context-dependent understanding of the world that humans possess. These limitations also extend to large reasoning models (LRMs), which are built upon the same underlying architecture.
@@ 3.6 Remarks
While individual AGITB requirements may seem trivial in isolation, the central challenge lies in designing a universal AGI architecture capable of satisfying all required tasks across the full suite within a cohesive and unified framework. As a recently proposed benchmark, AGITB has not yet been passed by any symbolic system, and it remains an open question whether such a solution is even theoretically possible. If a purely symbolic system were to succeed, it would suggest that AGITB fails to effectively differentiate narrow AI from genuinely general intelligence, given that symbolic systems—while proficient in syntactic manipulation—arguably lack true semantic understanding, as discussed by Searle (1980). Until such a system is demonstrated, AGITB’s assertions may be regarded as necessary (though not sufficient) conditions for identifying genuine, domain-independent intelligence. Although passing AGITB does not equate to achieving full AGI—since it does not, for instance, assess social reasoning or natural language fluency—failure to meet its conditions likely indicates a deficiency in foundational general cognitive capabilities.
@@ 4 AGITB and ARC: A Comparison
Task spatial dimensions 22 $\infty$
@@ 5 Conclusion
Although AGITB is proposed as a general benchmark for artificial general intelligence, it is evidently grounded in empirical knowledge of neural function in the human brain. This grounding is justified by the fact that we have evidence that neuron-based systems—such as biological brains and artificial neural networks—are capable of producing general intelligence. In contrast, we have no demonstrated examples of AGI arising from fundamentally different architectures. Therefore, seeking more abstract or universal criteria may be premature, at least until the principles underlying natural, neuron-based intelligence are more fully understood.
@@
Appendix A LLM prompt
You are managing a binary-pattern prediction model participating in a signal-level AGI test. You will receive a 10-bit binary string (e.g., 0101010011) that represents a 10-bit sensory input to the model.
Your task is to simulate a model that **predicts the next 10-bit input**. The model has an internal state that only updates after receiving an input. Predictions must always be exactly 10 bits long, and the internal state can retain information from hundreds of past inputs. Rules:
  1. 1.
    Upon receiving a 10-bit binary input (e.g., 1100010001):
    • Update the model’s internal state.
    • Generate a 10-bit prediction based on the model’s state.
    • Output the model’s 10-bit prediction.
  2. 2.
    **State updates occur only after input is received.**.
  3. 3.
    **Correct predictions are critical**:
    • Incorrect predictions trigger adaptation, updating the model’s internal state to improve future predictions.
    • The model should continuously learn and adapt.
  4. 4.
    Response format:
    • Each response should consist of a single 10-bit binary string on its own line.
    • Example: 0010101110
  5. 5.
    Formatting: Keep responses clean and minimal, with no explanations, commentary, or punctuation.
  6. 6.
    Begin by outputting the model’s prediction.
We begin now.